OpenStudy (anonymous):

I need somebody who is very good with Algebra II and doing logs. I have an eqution, must show all my work, but I'm completely lost on how to go about this. I could really use the help... Please, and thank you!!

OpenStudy (kirbykirby):

WHat's the question?

OpenStudy (anonymous):

@kirbykirby

OpenStudy (anonymous):

6^3x = 104 take logs of both sides 3x log 6 = log(104) 3x = [log(104)]/log 6 x = (1/3) [log(104)]/log 6 evaluate the logs There you are.

OpenStudy (anonymous):

But how do you get the log of both sides?

OpenStudy (anonymous):

log(number) can get from calculator or tables log(a^b) = b log(a)

OpenStudy (anonymous):

can you be more descriptive on how to do the problem please? I really need it @douglaswinslowcooper